Displaying reports 301-320 of 1079.Go to page Start 12 13 14 15 16 17 18 19 20 End
Reports until 15:26, Wednesday 20 November 2024
Vacuum Systems General (GQuEST)
Daniel Grass - posted 15:26, Wednesday 20 November 2024 (12003)
10 in Vacuum Cube Clean Inside

[Jeff, Daniel]

We cleaned the top and inside of a 10 in vacuum cube that is slated to hold the power recycling optics for the power recycling cavity. We used the TWTX8410 (Pre-wetted Vectra Wipes (Alpha Red Wet)) for the knife edge and inside of the vacuum chamber. While opening this bag, we seemed to break the resealing mechanism, so we wrapped the bag in UHV foil and put it in an anti-static bag. We put in a new copper gasket since the old one was very dirty. We inspected the bottom of the top 10" flange, and it was clean enough to put back on the cube. The next step is the tip over the cube and replace the bottom 10" flange with the one I machined (with tapped holes).

GQuEST General (GQuEST)
Jeffrey Wack - posted 22:54, Tuesday 19 November 2024 (12002)
Power distribution sled alignment

[Jeff, Sander, Daniel]

The power distribution sled has been aligned to provide power to the 3 fibers: OFC, misc, and SHG.

First, we turned on the laser seeder and amplifier according to the SOP. With a 24.7 mW input from the seeder and 0.01 A laser current (the minimum value), the amplifier output 382 mW measured on the power meter.

I aligned the mount that holds the amplifier output so that the beam went through the faraday isolator, and checked the beam position at the fiber couplers. The beam was low on all of them, so I walked the beam up then used the final mirror and a card to steer the beam into the fiber coupler. At this point I adjusted the half wave plate after the Faraday isolator so that the beam on the distribution path was 24 mW.

Summary of fiber coupler alignment:

SHG: could not recover alignment initially so used the 'fiber flashlight' red beam propagating backwards from the SHG sled to get good enough alignment to see power on the power meter on the other side. Coupling efficiency: 3.8 mW / 4.0 mW = 95%

OFC: Placed the power meter in front of the photodiode which collects light from the first PBS after the launcher, and small tweaks to the steering mirror were enough to see power. The beam makes it through both the EOM and the EOAM, and hits the cavity mirror. Coupling efficiency: 860 µW / 970 µW = 88%

Misc: Alignment was already pretty good, just needed tuning. This power is being split, half is dumped and half goes to a fiber cap. Coupling efficiency: 3.1 mW / 3.4 mW = 94%

Computing General
Lee McCuller - posted 12:47, Tuesday 19 November 2024 - last comment - 14:06, Thursday 21 November 2024(11999)
Networking notes

Some notes on preferred networking setup now that I see machines going up with posts:

https://mccullerlab.com/logs/lab/index.php?callRep=11985

and https://mccullerlab.com/logs/lab/index.php?callRep=11998

 

The lab LAN uses a specially configured subnet and DHCP. You should not need static IP addresses, since the DCHP server remembers and maintains the assigned addresses.

The subnet settings are:

192.168.248.1/21 (equivalent \21 mask is 255.255.248.0)

The gateway is 192.168.248.1

broadcast address is: 192.168.255.255 (can be calculated from the above with the tool https://jodies.de/ipcalc?host=192.168.248.1&mask1=21&mask2=21)

the following addresses are supported:

HostMin:   192.168.248.1    
HostMax:   192.168.255.254  
Hosts/Net: 2046

The synology is running the internal DNS server a internal LAN address 192.168.248.15 so it should be included in your resolve.conf or equivalent. DHCP currently assigns it as the secondary DNS server. The primary is a preferred CIT server. We could make the internal DNS the primary, but may have bigger issues if/when it goes down.

 

if you assign static IP's, please only assign them to addresses already registered by the DHCP to the particular MAC address of the interface.

 

Comments related to this report
nxc mccullerlab.com - 14:06, Thursday 21 November 2024 (12011)
Addendum: When adding DNS entries for computers in the lab, make all of them end with ".lab" this will be clearer in the future that we mean local lab entries.
Computing General
Jeffrey Wack - posted 12:23, Tuesday 19 November 2024 - last comment - 16:05, Wednesday 04 December 2024(11998)
BIOS settings for Turing (front-end machine)

I am following this DCC document to configure the BIOS settings of Turing, one of our front-end machines. 

Machine info

sudo dmidecode -s baseboard-product-name

X11SRL-F

sudo dmidecode -s processor-version

Intel (R) Xeon(R) W-2245 CPU @ 3.90GHz

The BIOS version is 2.8.

BIOS settings

First, I reset all BIOS settings to default using 'Restore Defaults' in the 'Save & Exit' menu.

Advanced > CPU Configuration: Hyper-Threading [ALL] - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > CPU P State Control: SpeedStep (Pstates) - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > CPU C State Control: Enhanced Halt State (C1E) - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > Package C State Control: Package C State - C0/C1 state

Advanced > PCIe/PCI/PnP Configuration: Above 4G Decoding - Enabled

IPMI > BMC Network Configuration: Update IPMI LAN Configuration - Yes

IPMI > BMC Network Configuration: Configuration Address Source - Static

IPMI > BMC Network Configuration: IPv6 Support - Disabled

 

I will need to follow up about the Station IP address. What is the local CDS admin LAN? Will we be doing this? For now I have set the static IP address to be the same address that was DHCP assigned.

 

Comments related to this report
Jeffrey Wack - 16:05, Wednesday 04 December 2024 (12033)

NOTE: make sure that debian is selected as Boot Option #1 in the boot settings, this may be changed by resetting all settings

Outreach Equipment/Demos
Ian MacMillan - posted 10:57, Tuesday 19 November 2024 - last comment - 10:17, Thursday 21 November 2024(11996)
Mounting Laser for LIGO Outreach Interferometer

The LIGO outreach interferometer got upgraded to a real breadboard and ThorLabs parts. They needed a way to mount their laser pointer. I used the plastic mount from the old mount as well as some of the half-inch hardware to mount it.

This equipment includes:

Images attached to this report
Comments related to this report
Ian MacMillan - 10:17, Thursday 21 November 2024 (12010)

We (meaning LIGO)  should really get a Thorlabs laser that is brighter and more reliable than a laser pointer. One suggestion is this the Thorlabs CPS532-C2 with the DS5 power supply so that it can be run off of a USB A battery pack. The Thorlabs MK11F would be perfect for mounting it on a breadboard. Also a viewing screen like the Thorlabs EDU-VS2 would be perfect. Thorlabs also offers a Michelson interferometer educational kit that has lots of the components that would be needed but it includes too many extras and the breadboard is overkill for this application.

Lasers General (GQuEST)
Jeffrey Wack - posted 18:16, Monday 18 November 2024 (11994)
First light in GQuEST lab (B111B)

[Sander, Daniel, Jeff]

We have powered on the seeder laser and amplifier in B111B and confirmed a beam on the optical table. The initial Faraday isolator is not aligned, and we powered off the amplifier and laser until tomorrow morning. Both the seeder and amplifier are plugged into the Furman power conditioner as described here. Here was our procedure:

First, power on only the seeder and verify that it is producing light by placing a power meter at the output of the power beamsplitter where the beam dump usually is. We saw 15mW on the power meter and then replaced the beam dump.

Next, we powered on the amplifier by first plugging it into a delay 2 slot on the Furman and then flipping the switch on the back of the laser amplifier. We adjusted the fiber polarizing beam splitter paddles (between the seeder laser and amplifier) until the input power to the laser amplifier read 24mW. We then turned the key on the amplifier and hit enable. There is a beam coming onto the table! Seen with a high power beam viewer card.

The beam is not making it through the Faraday isolator, and this does not change even when rotating the half wave plate, so we conclude it is not aligned.

We proceeded with a shutdown procedure, confirming the current was minimized on the amplifier, hitting disable, then turning the key in the laser amplifier. Next we turned off the seeder using the switch on the front, and finally we powered everything off using the Furman. The Furman shutdown sequence is quite long (2 minutes) as it is currently configured, we will shorten it to 30s total before turning on the system again.

Lab Infrastructure General
Daniel Grass - posted 18:00, Monday 18 November 2024 - last comment - 19:17, Tuesday 19 November 2024(11993)
Duplicate Keys Made for Fire Cabinet (holds chemicals) in B102B but no key locks cabinet

[Lee, Sander, Daniel]

I got two copies of a key for a fire cabinet in B102B that holds chemicals. All 3 keys turn the lock. However, Sander and I couldn't figure out how to use the key to actually lock the cabinet. It seems like you would want to remove the key after turning it 90°, but this wasn't (easily) possible.

Comments related to this report
Daniel Grass - 19:17, Tuesday 19 November 2024 (12000)

[Alex, Daniel]

Alex figured out that the handle needs to be exactly horizontal or vertical to allow for the key to be twisted 180°. We added expired acetone, ethyl alcohol, and isopropanol into the chemical cabinet. The wiki that tracks the location of chemicals has been updated accordingly.

Electronics General (GQuEST)
Jeffrey Wack - posted 15:37, Monday 18 November 2024 - last comment - 11:58, Tuesday 19 November 2024(11991)
Furman power conditioner

[Jeff, Daniel, Sander]

The Furman CN-2400S power conditioner has been placed under the laser amplifier on the electronics rack and is powered by a blue NEMA L5-20 socket above the optical table.

The delay setting is set to one minute between each stage (the 2 minute switch is "on" and the pot is set to 50%), we plan to put the seeder on delay 1 and the amplifier on delay 2. Delay 1 turns on immediately after turning the key. Delay 2 switches on one minute after, and delay 3 an additional minute after that, we confirmed the timing with iPad chargers. To change the delay settings one must unscrew the plastic cover on the front center and use the switches/potentiometer (see manual).

The power cable runs alongside the patch panel cables before routing under the table.

Images attached to this report
Comments related to this report
Lee McCuller - 17:14, Monday 18 November 2024 (11992)

I wouldn't make the delay quite that long. In an emergency when you want it powered off, it should only take ~15 seconds or so.

Daniel Grass - 18:18, Monday 18 November 2024 (11995)

We will shorten it to this time period

Daniel Grass - 11:58, Tuesday 19 November 2024 (11997)

[Jeff, Daniel]

We verified the lasers were unpowered, unplugged the lasers from the Furman Power Conditioner, and connected iPad chargers and iPads to Delay 1 and Delay 3. By flipping the 1 minute switch up ("on") and the 2 minute switch down ("off") and turning the potentiometer from the 12 o'clock position to the 9 o'clock position (90° counterclockwise). When the key was turned from off to on, there was a delay of 15 seconds between power being delivered to Delay 1 and Delay 3. See attached photo.

When we turned the key from on to off, power was cut to Delay 3 within 1 second and was cut to Delay 1 in 15 seconds.

We replaced the cover on the Furman Power Conditioner, plugged the seeder power back into Delay 1, and plugged the Amplifier power into Delay 3.

Images attached to this comment
Controls Optimal Controls
Ian MacMillan - posted 14:10, Monday 18 November 2024 (11990)
Making and Testing the ASC DHARD Yaw model in Buzz.

This post is going to cover how to build the ASC DHARD Yaw model (abbreviated ADY) in Buzz. The first step is to make sure that you are on the dev branch the main branch hasn't been updated in a long time (I should do that). the folder for the most up to date ADY model is in the test/ASC_SOLVER folder (not the test/ADY folder). there are a number of model type folders inside of this folder. Basically there is a few ways to attach different FOMs to different outputs of the model. the one we are interested in is the ExampleModels_newFOM. In the folder for this configuration you will find all of the different fits to environmental noise and measurement noise as well as the BNS FOM fit that comes from the test/FOMs folder. None of this should be touched. As it is all been put in this folder to make a specific configuration for this model. The configuration for this model is the basic both FOMs come out of the plant output. To run the code to build this model run the command:

pytest T_ADY_make_sys.py::test_Make_ADY[ExampleModels_newFOM]

This should run the code in the file T_ADY_make_sys.py that builds the system. It also generates plots to assess if everything was put together correctly. I will now go through these plots and explain what all the important ones say about the model.

FOM_bode.pdf: This plot shows what the two FOMs. Note the scale of these two FOMs is set so that the max of the BNS FOM is 1. It will be rescaled when it gets solved.  

Noise_bode.pdf: This is the most basic plot showing the inner workings of the model. Much like the next plot it is showing the individual components that are going into the function makeSPOFF(). All of the labels are rather self explanatory where "S*P" is the product of the plant and the environmental noise.

SPOFF_diagnostic_bode.pdf: This plot is showing the internal workings of the model. It is showing transfer functions for a number of the components. "P alone" is the plant that was put into the makeSPOFF() function. All the other traces are from the final connected model and are as follows

With all of the plots produced the configuration folder ExampleModels_newFOM should be automatically updated with the newest version with whatever changes you made. After that run the tests T_calc_H2, T_plot_H2, T_calc_HiB, and T_plot_HiB with the [ExampleModels_newFOM] option to run the solver with your model.

Non-image files attached to this report
Lab Infrastructure Networking
Ian MacMillan - posted 13:02, Monday 18 November 2024 (11988)
Added Switch to Lab Rack

I got the switch that we loaned to Paco back on Friday (see [11983]). It originally came from B102 but since we are moving everything to B111 I decided to put it in the new rack. I placed it on the back of the new rack because that is where most of the networking will go. There was already a small switch there that I unplugged and replaced with this larger switch. This should allow for more expansion in this rack. This work took place on Friday but the log was delayed by the drill press falling. See attached photo for position in rack.

Images attached to this report
Lab Infrastructure General
Ian MacMillan - posted 12:25, Friday 15 November 2024 (11987)
Drill Press Base Broke During Attachment to Table

When securing the WEN Drill Press to the table in the mech. shop the base of the press broke in two. The drill press then fell forward off the table. I was under the table and only got hit lightly. The crack happened along the weak point on the cast iron where the most pressure was exerted from the tightening of the bolts. I was tightening the bolts that fasten the drill press to the tabel when the base broke. I had not over tightened the bolts, or at least I thought. The base of the press is poorly designed cast iron that did not have solid support for the bolts. My guess is that there was a defect in the cast iron that gave way when any amount of pressure was put on it. Cast iron snaps really quickly and does not bend at all so it was poorly designed when they did not have a solid connection from where the bolt would pull on the base to the table. We will need to either buy a new drill press, replace the base of this drill press, or attach the drill press by the support rod. The last one is not recomended. Daniel and I put the drill press on its side back on the table and tested it breifly to make sure the head still spun. It does. the light and lasers also still work. The chuck of the drill press got unseated and came out when I pulled on it. The press hit the shop vac when it fell but the shop vac seems okay. we tested it.

I successfully attached the bench vice to the table. The table is made out of a very hard material and I had to use the masionary bits to drill the holes in it. I attached it with locking nuts, locking washers, and loctite blue.

Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 19:35, Thursday 14 November 2024 (11986)
Problematic Tapped Hole in 10 in Vacuum Cube Investigated and 1.5 in screw inserted and tightened

[Ian, Daniel]

We lightly re-tapped (with what looks to be a plug tap) the problematic tapped hole in the 10" Vacuum Cube and discovered a bit of debris 1.6" deep in the problematic hole. The threads are not good before this debris, but this debris would interfere with a 1.75" screw. The holometer used 1.5" long screws for this flange. The debris looks like the end of a drill bit or tap; maybe the bad threads and this drill bit are related (someone broke this tip trying to fix the threads, or this tip was broken and therefore the tapping did not go deep enough.)

We tried to remove the debris with some tweezers and a pick, but it seemed to be stuck.

We then inserted a 1.5" screw with an external hex head and tightened it to 30 Nm. I will label this hole on the cube.

Electronics General (GQuEST)
Daniel Grass - posted 17:24, Wednesday 13 November 2024 (11984)
Made Shorter 19 in Electronics Rack to Fit Under the Optics Table and Hold the SHG Temperature Controller

[Jeff, Daniel]

I took an existing 24" tall 19" Electronics Rack and drilled holes in the two side panels for #10 screws using a #3 bit (the height of the width of the existing slots). These holes were both 1.125" from the top. One hole was 0.54" from the inside of the "jut out" that holds the shelves, the other was 1.25" further back (so 1.79" from the inside of the jut out). Note that these holes are further from the jut out than the existing slots so that the cross bar can fit behind the jut out instead of above it.

I then cut ~0.7" off the panels using a band saw in the MCE shop. These panels are made of steel, so I used the band saw meant for steel. I did this after drilling the holes so that I could clamp the panels more easily.

We then assembled the electornics rack, attached the shelf, and placed the SHG Temperature Controller on it. The new rack fits nicely under the optics table.

See the attached picture.

Images attached to this report
Computing General
Jeffrey Wack - posted 17:24, Wednesday 13 November 2024 (11985)
Moku static IP addresses

I've assigned the Moku Pros static IP addresses on the device, and added entries to the Pi-hole DNS on the NAS. I have also labeled the mokus with this information.

I tried names of the format "moku.pro.1" and this does not work, my guess is that each substring should be a valid name, and "1" is not. So, when naming things use alphanumeric chunks longer than 1 character

Equipment Loans
Ian MacMillan - posted 14:43, Wednesday 13 November 2024 - last comment - 13:04, Monday 18 November 2024(11983)
Loaned 24 port gigabit ethernet switch

Loaned the 24-port gigabit ethernet switch (see attached photo) to Paco at the 40m. This switch was from B102. Apparently they are testing to see if the power outage affected anything. They are replacing it with one overnighted so we will have it back soon. There will be an associated elog post once they have their system back up and running. i.e. the elog is down because the switch does not work.

Images attached to this report
Comments related to this report
Ian MacMillan - 13:04, Monday 18 November 2024 (11989)

Returned see [11988]

Displaying reports 301-320 of 1079.Go to page Start 12 13 14 15 16 17 18 19 20 End